OpenStudy (anonymous):

What are the solutions of 6x^2 + 7x - 5 = 0?

OpenStudy (nathan917):

Simplifying 6x2 + 7x + -5 = 0 Reorder the terms: -5 + 7x + 6x2 = 0 Solving -5 + 7x + 6x2 = 0 Solving for variable 'x'. Factor a trinomial. (-5 + -3x)(1 + -2x) = 0 Subproblem 1 Set the factor '(-5 + -3x)' equal to zero and attempt to solve: Simplifying -5 + -3x = 0 Solving -5 + -3x = 0 Move all terms containing x to the left, all other terms to the right. Add '5' to each side of the equation. -5 + 5 + -3x = 0 + 5 Combine like terms: -5 + 5 = 0 0 + -3x = 0 + 5 -3x = 0 + 5 Combine like terms: 0 + 5 = 5 -3x = 5 Divide each side by '-3'. x = -1.666666667 Simplifying x = -1.666666667 Subproblem 2 Set the factor '(1 + -2x)' equal to zero and attempt to solve: Simplifying 1 + -2x = 0 Solving 1 + -2x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -2x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -2x = 0 + -1 -2x = 0 + -1 Combine like terms: 0 + -1 = -1 -2x = -1 Divide each side by '-2'. x = 0.5 Simplifying x = 0.5 Solution x = {-1.666666667, 0.5}

OpenStudy (anonymous):

0.5 or -5over3

OpenStudy (anonymous):

thnx

OpenStudy (ryan123345):